### v1

@@ Abstract
Surface ablation measurements of glaciers are critical for understanding mass change over time. Mass-balance stakes are commonly used for localized measurements, with the exposed length typically measured manually at infrequent intervals. This paper presents the design and validation of new instrumentation that automates mass-balance stake readings, thus enabling continuous measurements with high temporal resolution. The instrumentation comprises readout electronics that are mounted on mass-balance stakes to measure wind-induced vibrations. The stake vibrational frequency depends sensitively on the exposed length, and changes in the measured frequency therefore probe glacier surface melt and accumulation. Initial instrumentation field tests conducted at Color Lake on Umingmat Nunaat (Axel Heiberg Island), Nunavut, demonstrate centimeter-level precision on length measurements. The instrumentation can be attached to existing mass-balance stakes and is low-cost ($\sim$\$50 USD) in comparison to many other systems that perform automated surface ablation measurements. The accessibility of this instrumentation opens new possibilities for localized, high temporal resolution measurements of glacier surface activity at any locations where mass balance stakes are deployed.
@@
Glaciers and ice caps, recognized indicators of a changing climate, have exhibited continued and often accelerated mass loss through the early 21st century (e.g., hugonnet). These changes have strong implications for sea level rise, and for communities and ecosystems that depend on the runoff (immerzeel). The glaciological mass balance method (e.g., ostrem) is one approach used to quantify annual changes within a glacier system, using an input-output approach that interpolates and extrapolates measurements of mass gain (primarily snow accumulation) and mass loss (primarily ice melt and iceberg calving) across the glacier to gain a glacier-wide mass change commonly reported in meters of water equivalent (cogley). Point measurements of accumulation are made from measurements of snow thickness and density in snow pits, while ice melt is most commonly determined from measuring the exposure of mass-balance stakes drilled into the ice. Measurements for mass balance are most generally made only once or twice per year, representing annual or seasonal summer and winter balances, respectively (cogley). While this observation frequency is sufficient for standardized mass balance reporting, there is an increasing demand for data with higher temporal resolution to support and improve mass balance projections and runoff models. For most glacierized regions in the world, including the Canadian Arctic, ice melt is the dominant control on annual mass balance (sharp). Therefore, continuous melt monitoring can serve as a valuable, real-time, indicator of mass balance conditions and can support timely runoff projections for downstream environments and communities.
@@ 1 BRACHIOSAURUS design
The Boxed Recorder Analyzing the Change in Height of Ice with On-Site Accelerometer and Ultrasonic Readers Utilizing Support (BRACHIOSAURUS, or BRACHI for short) consists of an electronics box that is attached to the top of a mass-balance stake, as shown in Figure 1a. A microcontroller receives the data from an accelerometer, a temperature sensor, and a depth sensor (Figure 1b). The hardware design files, control code, and analysis software are all open source and available online. Each of the BRACHI components are described in detail below.
@@ 2 Vibrational frequency and exposed stake length
Figure 2 shows the numerically computed frequency–length relationship for the first two vibrational modes of an example BRACHI-equipped stake, along with the analytic prediction from Equation 1. (The detailed stake parameters are given in §5.) A small discrepancy between the analytic and numerical solutions for $F_{1}$ is visible at $\lesssim$1 m and arises from the differences in assuming an extended (numerical solution) or point mass (analytical solution) for BRACHI. In the remainder of this work, numerical solutions including an extended mass will be used for computing lengths from vibrational frequency measurements.
@@ 3 Data acquisition and analysis methods
The BRACHI firmware is configurable to allow user-defined data collection schedules. For the tests presented in this paper, data are recorded for 120 seconds every hour. The three-axis accelerometer is sampled at 200 Hz, and to reduce data volume, the acceleration magnitude squared is computed and saved to the micro SD card. The acceleration data are processed offline to determine the stake length as a function of time. Processing of each 120-second timestream begins with dividing the data into 10 equal-length chunks. Analyzing these chunks enables identification of low-amplitude vibrational modes that may fluctuate above and below the noise floor. Each chunk is Fourier transformed and squared to obtain a power spectrum, with the frequencies halved to correct for the doubling introduced by using the motion’s magnitude. In this spectrum, the vibrational frequencies of the stake appear as narrow peaks.
Figure 3 shows all of the frequency peaks that are typically present, although for some measurements under different conditions, only a subset of these frequencies may be excited at detectable levels. The peaks in the power spectrum include the first and second vibrational modes ($F_{1}$ and $F_{2}$) and other spectral lines at $2F_{1}$, $2F_{2}$, and $F_{2}\pm F_{1}$. The peaks at $F_{2}\pm F_{1}$ arise from Fourier transforming the squared magnitude of the acceleration.
@@ 5.2 Temporal variation of measured length
The BRACHI design files, control code, and analysis code are available on GitHub: https://github.com/FelixStA52/BRACHIOSAURUS
@@ Appendix A Detailed calculations of vibrational frequency as a function of length
as given in stokey. Here, $\mu(x)$ is the linear density of the box as a function of vertical position $x$, and $\phi_{n}(x)$ is the mode shape of the stake for the $n^{\text{th}}$ mode. The mode shape describes the displacement of the stake from equilibrium; the full expression is omitted here for brevity but is available in erturk_inman. Since $\phi_{n}(x)$ and $M_{\text{eff}}$ depend on each other, solutions for each must be obtained iteratively. Starting with $M_{\text{eff}}=m_{\text{BRACHI}}$, a trial mode shape is determined and used to compute a new effective mass. This process is repeated until the effective mass converges to the desired precision. The $M_{\text{eff}}$ solution is used to solve for $\lambda_{n}$ in Equation 2. The eigenvalues are related to the oscillation frequencies through the expression

### v2

@@ Abstract
Surface ablation measurements of glaciers are critical for understanding mass change over time. Mass-balance stakes are commonly used for localized measurements, with the exposed length typically measured manually at infrequent intervals. This paper presents the design and validation of new instrumentation that automates mass-balance stake readings, thus enabling continuous measurements with high temporal resolution. The instrumentation comprises readout electronics that are mounted on mass-balance stakes to measure wind-induced vibrations. The stake vibrational frequency depends sensitively on the exposed length, and changes in the measured frequency therefore are indicative of glacier surface melt and accumulation. Initial instrumentation field tests conducted at Color Lake on Umingmat Nunaat (Axel Heiberg Island), Nunavut, demonstrate centimeter-level precision on length measurements. The instrumentation can be attached to existing mass-balance stakes and is low-cost ($\sim$\$50 USD) in comparison to many other systems that perform automated surface ablation measurements. The accessibility of this instrumentation opens new possibilities for localized, high temporal resolution measurements of glacier surface activity at any locations where mass balance stakes are deployed.
@@
Glaciers and ice caps, recognized indicators of a changing climate, have exhibited continued and often accelerated mass loss through the early 21st century (e.g., hugonnet). These changes have strong implications for sea level rise, and for communities and ecosystems impacted by runoff (immerzeel). The glaciological mass balance method (e.g., ostrem) is one approach used to quantify annual changes within a glacier system, using an input-output approach that interpolates and extrapolates measurements of mass gain (primarily snow accumulation) and mass loss (primarily ice melt and iceberg calving) across the glacier to gain a glacier-wide mass change commonly reported in meters of water equivalent (cogley). Point measurements of accumulation are made from measurements of snow thickness and density in snow pits, while ice melt is most commonly determined from measuring the exposure of mass-balance stakes drilled into the ice. Measurements for mass balance are most generally made only once or twice per year, representing annual or seasonal summer and winter balances, respectively (cogley). While this observation frequency is sufficient for standardized mass balance reporting, there is an increasing demand for data with higher temporal resolution to support and improve mass balance projections and runoff models. For most glacierized regions in the world, including the Canadian Arctic, ice melt is the dominant control on annual mass balance (sharp). Therefore, continuous melt monitoring can serve as a valuable, real-time, indicator of mass balance conditions and can support timely runoff projections for downstream environments and communities.
@@ 1 BRACHIOSAURUS design
The Boxed Recorder Analyzing the Change in Height of Ice with On-Site Accelerometer and Ultrasonic Readers Utilizing Support (BRACHIOSAURUS, or BRACHI for short) consists of an electronics box that is attached to the top of a mass-balance stake, as shown in Figure 1 (a). A microcontroller receives the data from an accelerometer, a temperature sensor, and a depth sensor (Figure 1 (b)). The hardware design files, control code, and analysis software are all open source and available online. Each of the BRACHI components are described in detail below.
@@ 2 Vibrational frequency and exposed stake length
Figure 2 shows the numerically computed frequency–length relationship for the first two vibrational modes of an example BRACHI-equipped stake, along with the analytic prediction from Eq. (1). (The detailed stake parameters are given in §5.) A small discrepancy between the analytic and numerical solutions for $F_{1}$ is visible at $\lesssim$1 m and arises from the differences in assuming an extended (numerical solution) or point mass (analytical solution) for BRACHI. In the remainder of this work, numerical solutions including an extended mass will be used for computing lengths from vibrational frequency measurements.
@@ 3 Data acquisition and analysis methods
The BRACHI firmware is configurable to allow user-defined data collection schedules. For the tests presented in this paper, data are recorded for 120 seconds every hour. The three-axis accelerometer is sampled at 200 Hz, and to reduce data volume, the acceleration is squared and summed across all three axes before being saved to the micro SD card. The acceleration data are processed offline to determine the stake length as a function of time. Processing of each 120-second timestream begins with dividing the data into 10 equal-length chunks. Analyzing these chunks enables identification of low-amplitude vibrational modes that may fluctuate above and below the noise floor. Each chunk is Fourier transformed and squared to obtain a power spectrum, where the vibrational frequencies of the stake appear as narrow peaks.
Figure 3 shows all of the frequency peaks that are typically present, although for some measurements under different conditions, only a subset of these frequencies may be excited at detectable levels. The peaks in the power spectrum include the first and second vibrational modes ($F_{1}$ and $F_{2}$) and other spectral lines at $2F_{1}$, $2F_{2}$, and $F_{2}\pm F_{1}$. The peaks at $2F_{1}$, $2F_{2}$, and $F_{2}\pm F_{1}$ arise from Fourier transforming the squared magnitude of the acceleration. Future measurements will save and process the motion of all axes separately without squaring.
@@ 5.2 Temporal variation of measured length
The BRACHI design files, control code, analysis code, and data gathered at MARS and in the lab are available on GitHub: https://github.com/FelixStA52/BRACHIOSAURUS or under the following DOI: 10.5281/zenodo.18292185.
@@ Appendix A Detailed calculations of vibrational frequency as a function of length
as given in stokey. Here, $\mu(x)$ is the linear density of the box as a function of vertical position $x$, and $\phi_{n}(x)$ is the mode shape of the stake for the $n^{\text{th}}$ mode. The mode shape describes the displacement of the stake from equilibrium; the full expression is omitted here for brevity but is available in erturk_inman. Since $\phi_{n}(x)$ and $M_{\text{eff}}$ depend on each other, solutions for each must be obtained iteratively. Starting with $M_{\text{eff}}=m_{\text{BRACHI}}$, a trial mode shape is determined and used to compute a new effective mass. This process is repeated until the effective mass converges to the desired precision. The $M_{\text{eff}}$ solution is used to solve for $\lambda_{n}$ in Eq. (2). The eigenvalues are related to the oscillation frequencies through the expression